# Differentially Private Low-Rank Adaptation of Large Language Model Using Federated Learning

### v1

@@ 3. Background of Differential Privacy
Distributed training. We consider distributed training of a deep neural network with parameter $\bm{\theta}$. There are $K$ nodes participating in distributed training. Each node $k$ performs the gradient descent using its private dataset $D_{k}$, which contains $N_{k}$ samples. Thus, for node $k$, we have the sampling probability $q_{k}=B/N_{k}$. There are $N$ samples in total, where $N=\sum_{k=1}^{K}N_{k}$. In particular, one distributed training iteration consists of the following three:
@@
Theorem 1 ().
(Sequential composition (McSherry, 2009)) Let $\mathcal{M}_{k}$ each provides $(\epsilon_{k},\delta_{k})$-differential privacy, $k=1,2,...,K$. The sequential composition of $\mathcal{M}_{k}$ provides $(\epsilon_{k}^{{}^{\prime}},\delta_{k}^{{}^{\prime}})$-differential privacy, where the $\epsilon_{k}^{{}^{\prime}}=\sum_{k}\epsilon_{k}$ and $\delta_{k}^{{}^{\prime}}=\sum_{k}\delta_{k}$.
@@
Figure 1. Differential private low-rank adaptation scheme for LLMs using federated learning. Each node communicates with the server in updating $\bm{\theta}$.
@@ 4. Differentially Private Low-Rank Adaptation (DP-LoRA) of LLMs
Current advancements in large language models have led to an increased focus on open-source models by various research institutions, such as the Linux Foundation, Hugging Face, and AI4Finance Foundation. These organizations aim to foster large language models in vertical domains through collaborative efforts. For instance, banks are coming together to train financial models on transaction strategies and risk management and hospitals collaborate to develop models for diagnosing and treating medical conditions. These collaborations often involve fine-tuning existing large language models with private data to adapt them to specific tasks.
Each participating institution in such collaborations possesses sensitive, non-public data. Due to the constraints of privacy laws, this data cannot be shared among collaborating entities. As a result, institutions adopt the federated learning approach, in which each participant trains large language model locally and shares only the model weights in collaboration. However, studies have discovered that it is possible to extract training data from the weights of large language models. This data extraction attack means that the current method of sharing model weights in collaborations still poses a risk of private data leakage. Given these risks, there is an urgent need to establish a mechanism that provides privacy protection in collaborative model development.
We propose a differentially private low-rank adaptation (DP-LoRA) method of LLMs in federated learning. We adopt a Gaussian mechanism that adds noise to the model weights during the training process. Our method introduces a privacy guarantee for low-rank adaptation which can also reduce the communication overhead in federated learning. First, we discuss differential private low-rank adaptation in each node. We then address the communication scheme between each node and the server.
4.1. Problem Formulation
In this subsection, we formally describe the problem of differentially private distributed training for LLMs. In what follows, we first introduce differentially private distributed training and analyze the communication overhead. Then we describe our objective formally.
Differentially private distributed training. DP-SGD can also generalize to distributed training. Each node $k$ adopts (6) to calculate gradient $\widetilde{\bm{g}_{k}}$ and updates $\bm{\theta}_{k}$ by (3). Therefore, in node $k$, we can guarantee the ($\epsilon_{k},\delta_{k}$)-differential privacy for one iteration by adding Gaussian noise with mean 0 and standard deviation $\sigma\geq C\sqrt{2\ln(1.25/\delta_{k})}/\epsilon_{k})$.
Theorem 1 ().
(Parallel composition (McSherry, 2009)) Let $\mathcal{D}$ be a domain, let $D\in\mathbb{N}^{\mathcal{D}}$ be a data set, and let $K$ be a positive integer. For $k=1,2,...,K$, let $D_{k}$ be a subset of $\mathcal{D}$, let $\mathcal{M}_{k}$ be a mechanism that takes $D\cap D_{k}$ as input, and suppose $\mathcal{M}_{k}$ is $(\epsilon_{k},\delta_{k})$-differentially private. If $D_{i}\cap D_{j}=\emptyset$ whenever $i\neq j$, then the composition of the sequence $\mathcal{M}_{1},\ldots,\mathcal{M}_{k}$ is $(\max\{\epsilon_{i}\},\max\{\delta_{i}\})$-differentially private, where $i=1,...,k$.
As each $D_{k}$ came from a different distribution, the composition of $K$ differential private mechanism satisfies $(\max(\epsilon_{k}),\max(\delta_{k}))$-differential private. Therefore, one iteration of global model updates is a differentially private mechanism. The whole training process conducts $T$ iterations. Following Theorem 7, we can have the noise scale to guarantee the whole training process as a differential private mechanism, which can be represented as $\Omega(\frac{\max(q_{k})\sqrt{T\log(1/\delta)}}{\epsilon})$, where $q_{k}=B/N_{k}$.
Communication overhead. The communication overhead in federated learning involves several factors that contribute to the amount of data transmitted and the efficiency of this transmission. The communication overhead in federated learning is mainly related to five elements, which are model size, number of iterations, number of participants, and network conditions. The number of participants and update frequency can be adjusted easily and we assume the network condition is ideal without any noise. Thus, in this paper, we assume the communication overhead is mainly determined by the transferring model size. Unfortunately, LLMs often have a large number of parameters, which incurs significant communication overhead with federated learning.
Differentially private distributed training for LLMs. Given $K$ parties (e.g., hospitals or banks), where each party owns a private text dataset $D_{k}$ ($k=1,2,...,K$) for fine-tuning an LLM. We aim to develop a differentially private distributed training algorithm for the LLM, such that (1) the ($\epsilon_{k},\delta_{k}$)-differential privacy can be guaranteed, and (2) the communication overhead is minimized.
4.2. Proposed DP-LoRA Algorithm
@@
Fig. 1 shows the architecture of our proposed DP-LoRA algorithm. Every node $k$ has a local dataset $D_{k}$. The dataset $D_{k}$ is not shared with the server or other nodes and has $N_{k}$ samples. We use $N$ to denote the total number of samples, where $N=\sum\limits_{k=1}^{K}N_{k}$. For the large language model, each node $k$ freeze the original model weight $\bm{W}$ and fine-tune the $\bm{\theta}_{k}$, where $\bm{\theta}_{k}$ is composed by low-rank decomposition matrix of each attention head. For each transformer layer $l$, the attention weight $\bm{W}^{l}$ is decomposed to $\bm{A}^{l}$ and $\bm{B}^{l}$. Therefore, $\bm{\theta}=\{\bm{A}^{1},\bm{B}^{1};\bm{A}^{2},\bm{B}^{2};...;\bm{A}^{L},\bm{B}^{L}\}$.
@@
    Broadcast: The server sends the global model parameters $\bm{\theta}$ to all participating nodes. This is the starting point for training in each round.
@@
    Local Training: Each node $k$ trains the model locally on its own dataset $D_{k}$. The freeze $\bm{W}$ indicates that the original model only does inference during local training and the weights are not updated by the gradient. We obtain a new $\bm{\theta}_{k}$ using gradient descent.
@@
    Add Noise: A Gaussian noise $n_{k}\sim\mathcal{N}(0,\sigma^{2})$ is add to the $\bm{\theta}_{k}$ such that $\widetilde{\bm{\theta}_{k}}=\bm{\theta}_{k}+n_{k}$.
@@
    Upload: Every node $k$ uploads $\widetilde{\bm{\theta}_{k}}$ to the server.
@@
    Aggregate: The server aggregates the $\widetilde{\bm{\theta}_{k}}$ from all nodes and takes the weighted average $\sum\limits_{k=1}^{K}\rho_{k}\widetilde{\bm{\theta}_{k}}$ to updates $\bm{\theta}$, where $\rho_{k}=N_{k}/N$.
We elaborate step $2$ and step $3$ by Alg. 1, which can be exemplified as DP-LoRA. Given a local dataset $D_{k}$, a noise scale $\sigma$, a clipping norm $C$, a number of iterations $T$, a learning rate $\gamma$, and a batch size $B$, we initialize local model parameters $\theta_{k}$. The general process can be described as
  1. (1)
    Select a random batch: Choose $B$ samples from the dataset $D_{k}$ randomly.
  2. (2)
    Compute gradients: Calculate the gradients of the loss function $\mathcal{L}$ with respect to the parameters $\theta_{k}$, denoted as $\bm{g}_{k}$.
  3. (3)
    Clip gradients: Perform gradient clipping by scaling down the gradients if their L2 norm is greater than $C$, $\bm{g}_{k}\leftarrow\bm{g}_{k}/\max(1,\frac{\|\bm{g}_{k}\|_{2}}{C})$.
  4. (4)
    Add noise: Introduce Gaussian noise to the clipped gradients to ensure differential privacy, $\bm{g}_{k}\leftarrow\bm{g}_{k}+\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$, $k=1,2,...,K$.
  5. (5)
    Gradient descent update: Update the local model parameters via gradient descent, $\widetilde{\theta_{k}}\leftarrow\theta_{k}-\gamma{\bm{g}_{k}}$, where $\bm{g}_{k}$ is the noised and clipped gradient.
4.3. Guarantee of Locally Differential Privacy
We assume that there are $K$ nodes in total. Each node $k$ has a private dataset $D_{k},k=1,2,...,K$. Each node holds an initial weight $\bm{W}$ which is freezed and only used for making inferences. All nodes hold the same $\bm{W}$. The node $k$ uses the dataset $D_{k}$ in fine-tuning the low-rank adaptation modules with parameter $\bm{\theta}_{k}$.
We introduce a differentially private low-rank adaptation (DP-LoRA) as a local training algorithm. We add noise to the gradient in each iteration. For $T$ training iterations, the proposed algorithm can satisfy $(\epsilon,\delta)$-differential privacy. Our algorithm is presented as follows.
@@ 4.3. Guarantee of Locally Differential Privacy
5:     $\bm{\theta}_{k}\leftarrow$ Nodeupdate($D_{k}$, C, B, $\sigma$, $\gamma$),   $k=1,2,...,K$
6:     $\bm{\theta}\leftarrow\sum\limits_{k=1}^{K}\rho_{k}\bm{\theta}_{k}$
@@
11:  sample $(\bm{x},\bm{y})$ from $D_{k}$ with batch size $B$
12:  $\bm{g}_{k}\leftarrow\nabla_{\bm{\theta}_{k}}\mathcal{L}(\bm{\theta}_{k};\bm{x},\bm{y})$
13:  $\bm{g}_{k}\leftarrow\bm{g}_{k}/\max(1,\frac{\|\bm{g}_{\bm{\theta}_{k}}\|_{2}}{C})$
14:  ${\bm{g}}_{k}\leftarrow{\bm{g}}_{k}+\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$
15:  $\bm{\theta}_{k}\leftarrow\bm{\theta}_{k}-\gamma\bm{g}_{k}$
16:  Return $\bm{\theta}_{k}$
17:  end function
A local model is composed by $L$ weight matrix $\bm{W}^{l}\in\mathbb{R}^{n\times n},l=1,2,...,L$. Here, we consider the linear layers, and the input of output shares the same dimension $n$. These weight matrices have a low “intrinsic dimension" (Aghajanyan et al., 2020). We decompose the matrix $\bm{W}^{l}$ into $\bm{A}^{l}\in\mathbb{R}^{n\times r}$ and $\bm{B}^{l}\in\mathbb{R}^{r\times n}$, follow by low-rank adaptation method (Hu et al., 2021). Note that the rank $r\ll n$. Thus, we have $\bm{\theta}=\{\bm{A}^{1},\bm{B}^{1};\bm{A}^{2},\bm{B}^{2};...;\bm{A}^{L},\bm{B}^{L}\}$. We follow the setting in (Abadi et al., 2016) to sample $B$ data from the dataset to provide an unbiased estimation.
Thirdly, we add Gaussian noise in the gradient. The probability density function for Gaussian noise is $\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$, where $\bm{I}$ is an identity matrix and has the same shape with $\bm{\theta}_{k}$. The clipping bound $C$ is used to ensure the gradient norm $||\bm{g}_{k}||<C$. The noise scale $\sigma$ provides the privacy guarantee for each iteration, which should follow $\sigma\geq\sqrt{2\ln{(1.25/\delta)}}/\epsilon$. We update the add noise $\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$ into matrix $\bm{A}$ and $\bm{B}$, respectively.
Equation (6) guarantees that one training iteration is $(\epsilon,\delta)$-differential privacy. Alg. 1 composes $T$ iterations. As the privacy amplification theorem (Hu et al., 2021; Beimel et al., 2014) and strong composition theorem (Dwork et al., 2014), the composition of $T$ differential private mechanism is still differential private. The algorithm 1 is $(\epsilon,\delta)$-differential private.
@@ 4.4. Reducing Communication Overhead in Federate Learning
We assess the reduction in communication overhead achieved through the low-rank adaptation method. Assuming that the federated learning process spans $T$ rounds in a noiseless and ideal network environment, the communication overhead can be quantified using specific parameters. The proposed method provides the reduction in two steps. First, we freeze most of the model weight. We fine-tune and transfer the weight in self-attention layers. For example, a Llama-7B is composed of $32$ transformer layers and each transformer consists $32$ attention heads, which have $4096$ dimensions. Thus, each self-attention block contains $4096\times 4096\times 3=50,331,648$ parameters, and $32$ transformers contain $1,610,612,736$ parameters in total. Therefore, we reduce the training parameter from $6,738,411,520$ to $1,610,612,736$. Second, we decomposed the self-attention parameters by low-rank adaptation methods. We use a $\bm{W}^{l}$ to denote the parameters of one attention head. Each $\bm{W}^{l}\in\mathbb{R}^{n\times n}$ is replace by a $\bm{A}^{l}\in\mathbb{R}^{n\times r}$ and $\bm{B}^{l}\in\mathbb{R}^{n\times r}$. Thus, for one attention head with $4096$ dimension, we reduce the parameter from $16,777,216$ to $2,097,152$ if we choose $r$ as $256$.

### v2

@@ 3. Background of Differential Privacy
Federated learning. We consider distributed training of a deep neural network with parameter $\bm{\theta}$. There are $K$ nodes participating in distributed training. Each node $k$ performs the gradient descent using its private dataset $D_{k}$, which contains $N_{k}$ samples. Thus, for node $k$, we have the sampling probability $q_{k}=B/N_{k}$. There are $N$ samples in total, where $N=\sum_{k=1}^{K}N_{k}$. In particular, one distributed training iteration consists of the following three:
@@
Theorem 1.
(Sequential composition (McSherry, 2009)) Let $\mathcal{M}_{t}$ each provides $(\epsilon_{t},\delta_{t})$-differential privacy, $t=1,2,...,T$. The sequential composition of $\mathcal{M}_{t}$ provides $(\epsilon_{t}^{{}^{\prime}},\delta_{t}^{{}^{\prime}})$-differential privacy, where the $\epsilon_{t}^{{}^{\prime}}=\sum_{t}\epsilon_{t}$ and $\delta_{t}^{{}^{\prime}}=\sum_{t}\delta_{t}$.
@@
Figure 1. Differential private low-rank adaptation (DP-LoRA) algorithm for LLMs using federated learning. Each node communicates with the server in updating $\bm{\theta}$.
@@ 4. Differentially Private Low-Rank Adaptation (DP-LoRA) of LLMs
We propose a differentially private algorithm for a large language model with a federated learning scenario. First, we elaborate on the privacy guarantee in a non-balance federated learning of our methods. Second, our method is able to mitigate the communication overhead in LLM weight transmission.
4.1. Motivation
@@
Multiple institutes collaborate in training a large language model using their local datasets. Current advancements in large language models have led to an increased focus on open-source models by multiple institutions, such as the Linux Foundation, Hugging Face, and AI4Finance Foundation. These organizations collaborate in training large language models in vertical domains, such as medicine and finance. Each institute in the organization processes a local dataset for collaborative training. For instance, banks cooperate in training financial models on transaction strategies and risk management with their financial data. Hospitals collaborate to develop models for diagnosing and treating medical conditions with medical data. These collaborations often involve training existing large language models with data from the specific vertical domain. Each participant contributes their data in this collaboration framework.
Each participating institution usually possesses sensitive, private data. Due to the constraints of privacy laws, these data cannot be shared among collaborating entities. While federated learning has been widely studied for these scenarios, recent studies have discovered that the federated learning approach is not safe enough to protect private data in large language models (Carlini et al., 2021). In the meantime, the communication overhead becomes imminent as the model size becomes very large.
In this work, we proposed DP-LoRA, a cooperation scheme for large language model federated learning. Firstly, we prove that DP-LORA is a privacy-preserving mechanism that provides a privacy guarantee for sensitive training data. Secondly, we introduce the low-rank adaptation into federated learning to mitigate the communication overhead. Thirdly, we elaborate a weighted federate learning scheme that corresponds to varying data quantities and qualities.
4.2. DP-LoRA Algorithm
@@
Our method is shown in Fig. 1. There are $K$ nodes, where every node $k$ holds a local dataset $D_{k}$ and a large language model. For the dataset, $D_{k}$ has $N_{k}$ samples and is not shared with the server or other nodes. We use $N$ to denote the total number of samples, where $N=\sum\limits_{k=1}^{K}N_{k}$. For the large language model, every node has the same model weight $\bm{\theta}$ in the beginning. $\bm{\theta}$ includes the original LLM weight $\bm{W}$ and the low-rank adaptation weight $\bm{A}$ and $\bm{B}$. Each node $k$ freezes the original model weight $\bm{W}$ and fine-tunes the $\bm{A}_{k}$ and $\bm{B}_{k}$.
Given the noise scale $\sigma$, the clipping norm $C$, the number of iterations $T$, the learning rate $\gamma$, and the batch size $B$, we propose a differential private low-rank adaptation (DP-LoRA) algorithm for federated learning, while the pseudocode are summarized in Alg. 1.
@@
    Broadcast: in line 4 of Alg. 1, the server sends the global model parameters $\bm{\theta}$ to all participating nodes. This is the starting point for training in each round.
@@
    Local Training: In line 11-14 of Alg. 1, each node $k$ samples $B$ samples from $D_{k}$ and uses the clipped gradient to update the local weight $\bm{A}_{k}$ and $\bm{B}_{k}$.
@@
    Add Noise: In line 16 of Alg. 1, a Gaussian noise $n_{k}\sim\mathcal{N}(0,\sigma^{2})$ is add to the $\bm{A}_{k}$ and $\bm{B}_{k}$ simultaneously. We denote the noise weight as $\widetilde{\bm{A}_{k}}=\bm{A}_{k}+n_{k}$ and $\widetilde{\bm{B}_{k}}=\bm{B}_{k}+n_{k}$.
@@
    Upload: In line 5 of Alg. 1, every node $k$ uploads $\widetilde{\bm{A}_{k}}$ and $\widetilde{\bm{B}_{k}}$ to the server.
@@
    Aggregate: In line 6 of Alg.1, the server aggregates the $\widetilde{\bm{A}_{k}}$ and $\widetilde{\bm{B}_{k}}$ from all nodes and takes the weighted average $\sum\limits_{k=1}^{K}\rho_{k}\widetilde{\bm{A}_{k}}$ and $\sum\limits_{k=1}^{K}\rho_{k}\widetilde{\bm{B}_{k}}$ to updates $\bm{A}$ and $\bm{B}$ as follows
    (8) $\widetilde{\bm{A}}=\sum\limits_{k=1}^{K}(\bm{A}_{k}+n_{k})=\sum\limits_{k=1}^{K}\rho_{k}\bm{A}_{k}+\sum\limits_{k=1}^{K}\rho_{k}n_{k},\quad\widetilde{\bm{B}}=\sum\limits_{k=1}^{K}(\bm{B}_{k}+n_{k})=\sum\limits_{k=1}^{K}\rho_{k}\bm{B}_{k}+\sum\limits_{k=1}^{K}\rho_{k}n_{k}$
    where $\rho_{k}=N_{k}/N$.
We elaborate the process in lines 10-18 of Alg. 1. First, calculate the gradients of the loss function $\mathcal{L}$ for the parameters $\bm{A}_{k}$ and $\bm{B}_{k}$, denoted as $\bm{g}_{\bm{A}_{k}}$, $\bm{g}_{\bm{B}_{k}}$. Second, perform gradient clipping by scaling down the gradients if their L2 norm is greater than $C$, $\bm{g}_{\bm{A}_{k}}\leftarrow\bm{g}_{\bm{A}_{k}}/\max(1,\frac{\|\bm{g}_{\bm{A}_{k}}\|_{2}}{C})$ and $\bm{g}_{\bm{B}_{k}}\leftarrow\bm{g}_{\bm{B}_{k}}/\max(1,\frac{\|\bm{g}_{\bm{B}_{k}}\|_{2}}{C})$. Third, introduce Gaussian noise to the clipped gradients to ensure differential privacy, $\bm{g}_{\bm{A}_{k}}\leftarrow\bm{g}_{\bm{A}_{k}}+\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$ and $\bm{g}_{\bm{B}_{k}}\leftarrow\bm{g}_{\bm{B}_{k}}+\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$, $k=1,2,...,K$. Fourth, update the local model parameters via gradient descent, $\widetilde{\bm{A}_{k}}\leftarrow\bm{A}_{k}-\gamma\bm{g}_{\bm{A}_{k}}$ and $\widetilde{\bm{B}_{k}}\leftarrow\bm{B}_{k}-\gamma\bm{g}_{\bm{B}_{k}}$, where $\bm{g}_{k}$ is the noised and clipped gradient.
4.3. Guarantee of Differential Privacy
The differentially private low-rank adaptation (DP-LoRA) we introduced is a $(\epsilon,\delta)$ differential private mechanism. We prove that in Theorem 1.
@@ 4.3. Guarantee of Differential Privacy
5:     $\bm{A}_{k},\bm{B}_{k}$leftarrow Nodeupdate($D_{k}$, C, B, $\sigma$, $\gamma$),   $k=1,2,...,K$
6:     $\bm{A}\leftarrow\sum\limits_{k=1}^{K}\rho_{k}\bm{A}_{k}$,  $\bm{B}\leftarrow\sum\limits_{k=1}^{K}\rho_{k}\bm{B}_{k}$
@@
11:  Input: dataset $D_{k}$, clipping norm $C$, learning rate $\lambda$, batch size $B$ and noise scale $\sigma$.
12:  sample $(\bm{x},\bm{y})$ from $D_{k}$ with batch size $B$
13:  $\bm{g}_{\bm{A}_{k}}\leftarrow\nabla_{\bm{A}_{k}}\mathcal{L}(\bm{g}_{\bm{A}_{k}};\bm{x},\bm{y})$,  $\bm{g}_{\bm{B}_{k}}\leftarrow\nabla_{\bm{B}_{k}}\mathcal{L}(\bm{g}_{\bm{B}_{k}};\bm{x},\bm{y})$,
14:  $\bm{g}_{\bm{A}_{k}}\leftarrow\bm{g}_{\bm{A}_{k}}/\max(1,\frac{\|\bm{g}_{\bm{A}_{k}}\|_{2}}{C})$,  $\bm{g}_{\bm{B}_{k}}\leftarrow\bm{g}_{\bm{B}_{k}}/\max(1,\frac{\|\bm{g}_{\bm{B}_{k}}\|_{2}}{C})$
[…]
16:  $\bm{A}_{k}\leftarrow\bm{A}_{k}-\gamma\bm{g}_{\bm{A}_{k}}$,  $\bm{B}_{k}\leftarrow\bm{B}_{k}-\gamma\bm{g}_{\bm{B}_{k}}$
17:  Output $\bm{A}_{k}$, $\bm{B}_{k}$
18:  end function
For a single node $k$, we add Gaussian noise in the gradient when we update $\bm{A}_{k}$ and $\bm{B}_{k}$. The probability density function for Gaussian noise is $\mathcal{N}(0,\sigma^{2}C^{2}\bm{I})$, where $\bm{I}$ is an identity matrix and has the same shape with $\bm{A}_{k}$ and $\bm{B}_{k}$. The clipping bound $C$ is used to ensure the gradient norm $||\bm{g}_{\bm{A}_{k}}||<C$ and $||\bm{g}_{\bm{B}_{k}}||<C$. The noise scale $\sigma$ provides the privacy guarantee for each iteration, which should follow $\sigma\geq\sqrt{2\ln{(1.25/\delta)}}/\epsilon$. We use $\widetilde{\bm{A}_{k}}$ and $\widetilde{\bm{B}_{k}}$ to denote the noise weight from the node $k$.
For the server, it aggregates all $\widetilde{\bm{\theta}_{k}}$ uploaded from the server by Eq. (8). we prove that the differential private property holds for the proposed algorithm, which guarantees the privacy for the training dataset $D_{k}$.
Theorem 1.
There exists positive constants $c_{1}>0$ and $c_{2}>0$, the sampling probability $q=B/N$, the batch size $B$ and dataset size $N$, for any $\epsilon<c_{1}q^{2}T$, DP-LoRA with $K$ nodes in Alg. 1 over $T$ iterations is $(\epsilon,\delta)$-differentially private for a given $\delta>0$ such that
[…]
. In the above theorem, we assume that all the node adopts the same sample probability $q$. We detailed the proof in our appendix.
@@ 4.4. Reducing Communication Overhead in Federate Learning
In this section, we illustrate how the proposed DP-LoRA would reduce the communication overhead. Typically, a language model with wight $\bm{W}$ is composed by $L$ weight matrices $\bm{W}^{l}\in\mathbb{R}^{n\times n},l=1,2,...,L$. Here, we consider the linear layers, and the input of output shares the same dimension $n$. These weight matrices have a low “intrinsic dimension" (Aghajanyan et al., 2020). We decompose the matrix $\bm{W}^{l}$ into $\bm{A}^{l}\in\mathbb{R}^{n\times r}$ and $\bm{B}^{l}\in\mathbb{R}^{r\times n}$, follow by low-rank adaptation method (Hu et al., 2021). Note that the rank $r\ll n$. Thus, we have $\bm{\theta}=\{\bm{A}^{1},\bm{B}^{1};\bm{A}^{2},\bm{B}^{2};...;\bm{A}^{L},\bm{B}^{L}\}$.
@@
We assess the reduction in communication overhead achieved through the low-rank adaptation method. Assuming that the federated learning process spans $T$ rounds in a noiseless and ideal network environment, the communication overhead can be quantified using specific parameters. The proposed method provides the reduction in two steps. First, we freeze most of the model weight. We fine-tune and transfer the weight in self-attention layers. For example, a Llama-7B is composed of $32$ transformer layers and each transformer consists $32$ attention heads, which have $4096$ dimensions. Thus, each self-attention block contains $4096\times 4096\times 3=50,331,648$ parameters, and $32$ transformers contain $1,610,612,736$ parameters in total. Therefore, we reduce the training parameter from $6,738,411,520$ to $1,610,612,736$. Second, we decomposed the self-attention parameters by low-rank adaptation methods. We use a $\bm{W}^{l}$ to denote the parameters of one attention head. Each $\bm{W}^{l}\in\mathbb{R}^{n\times n}$ is replace by a $\bm{A}^{l}\in\mathbb{R}^{n\times r}$ and $\bm{B}^{l}\in\mathbb{R}^{n\times r}$. Thus, for one attention head with $4096$ dimension, we reduce the parameter from $16,777,216$ to $2,097,152$ if we choose $r=256$.